OpenStudy (anonymous):

What are the possible number of positive real, negative real, and complex zeros of f(x) = -7x4 - 12x3 + 9x2 - 17x + 3? Positive Real: 3 or 1 Negative Real: 1 Complex: 2 or 0 Positive Real: 3 or 1 Negative Real: 2 or 0 Complex: 1 Positive Real: 1 Negative Real: 3 or 1 Complex: 2 or 0 Positive Real: 4, 2 or 0 Negative Real: 1 Complex: 0 or 1 or 3

OpenStudy (jdoe0001):

have you covered Descartes Rule of Signs yet?

OpenStudy (anonymous):

nope, whats that?

OpenStudy (anonymous):

@Squirrels

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=5YAmwfT3Esc <--- Descarte rule of signs

OpenStudy (anonymous):

thank you :) but youtube is blocked on my computer ahaha oops but i apreciate your help thanks again

OpenStudy (anonymous):

if someone could help me i'd be really thankful!

OpenStudy (ranga):

f(x) = -7x^4 - 12x^3 + 9x^2 - 17x + 3 For Descartes Rule of Signs, assume x is positive. An easy thing to do would be to set x = +1 f(1) = -7 - 12 + 9 - 17 + 3 Count the number of times the terms change signs. one sign change from -12 to +9 then +9 to -17 then -17 to +3 Three sign changes. Therefore, there is a maximum of 3 positive roots. Try the same but assuming x is negative or set x = -1. Write down each term and see how many times the signs change. There will be a maximum of that many negative roots.

OpenStudy (anonymous):

Thank you :)

OpenStudy (ranga):

You are welcome. With positive x we got 3 sign changes. That means maximum is 3 positive. But it could also be just1 positive. You always step down 2 at a time.

OpenStudy (anonymous):

Okay, so the positive is 3 &1. What about the negative & complex

OpenStudy (ranga):

f(x) = -7x^4 - 12x^3 + 9x^2 - 17x + 3 Let us try x = -1 f(-1) = -7(-1)^4 - 12(-1)^3 + 9(-1)^2 - 17(-1) + 3 = -7 + 12 + 9 + 17 + 3 There is just ONE sign change from -7 to +12. The rest are all + so no more sign changes. Here it is definite. There is one negative root. (had there been say 2 sign changes here we will say, it can have 2 negative roots or 0 negative roots (always step down by 2 from the maximum possibilities)).

OpenStudy (anonymous):

Positive Real: 3 or 1 Negative Real: 1 Complex: 2 or 0 this is what i got.

OpenStudy (ranga):

This is a fourth degree polynomial. Therefore it will have a total of 4 roots. We know for sure that it has 1 negative root. That leaves with 3 roots. If positive is 3, complex must be 0. If positive is 1, complex must be 2. Yes, your answer is correct.